OpenStudy (anonymous):

Solve the following system of equations using Substition y= -2x - 2 and 9x + 4y = -11

OpenStudy (anonymous):

u can solve this well x first by plugging in -2x-2 into 4 so.. 9x+4(-2x-2)=-11 and solve for x! :)

OpenStudy (anonymous):

u dont get it?

OpenStudy (anonymous):

and thats using substitution?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

cuz u have y= so that means that whatever y= or x= u'll just plug it into the other equation. :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome! i hope i helped! :l

OpenStudy (anonymous):

so 9x+4(-2x-2)=-11 do i distrubute the 4 over?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

so x=3?

OpenStudy (anonymous):

let me check.....

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

its -3

OpenStudy (anonymous):

ok and that is all i need to do that will be the answer to the whole equation?

OpenStudy (anonymous):

no...... u need to solve for y now so basically plug in -3 into the first equation y= thats the easier one and ull get ur x :)

OpenStudy (anonymous):

so y=4?

OpenStudy (anonymous):

yes!!!! good job! :))

OpenStudy (anonymous):

XD Thanks May i ask you others?

OpenStudy (anonymous):

depends on how many and r they similar?? cuz though im taking pre-calc i'm a little rusty on some math problems

OpenStudy (anonymous):

ok they kinda are

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

Solve the follwoing system of equations ing Linear Combinations 4x + y and -3x - 5y = -8

OpenStudy (anonymous):

4x+y does not have an = sign?

OpenStudy (anonymous):

sorit is 4x + 2y = 20 and -3x - 5y = -8

OpenStudy (anonymous):

okay for this one u would solve for y again id choose 4x+2y=20 cuz it easier so u would solve for y... do u know how 2 do that?

OpenStudy (anonymous):

no plz help

OpenStudy (anonymous):

okay so u have 4x+2y=20 u would divide by 2 on both sides u would have 2x+y=10 then u'll move the 2x on the other side so u'll have y=-2x+10 now u can plug y into the other equation get it????

OpenStudy (anonymous):

kinda would i do -3x - 5(-2x + 10)=-8

OpenStudy (anonymous):

which equation did u plug it into?

OpenStudy (anonymous):

-3x - 5y = -8

OpenStudy (anonymous):

wasn't it 9x+4y=-11?

OpenStudy (anonymous):

Solve the follwoing system of equations ing Linear Combinations 4x + y and -3x - 5y = -8 this is the new equation

OpenStudy (anonymous):

oooo u shouda gave me head up then :l ahhh. okay so does 4x+y have an equal sign? wait didnt we do this one already???

OpenStudy (anonymous):

no we didnt in my packet question number 14 is

OpenStudy (anonymous):

okay NEVERMIND what u said before about plugging it into the-3x-5y=-8 is right sorry im like losing track

OpenStudy (anonymous):

its ok want me to stop?

OpenStudy (anonymous):

no its fine for now :) is it fine if i let u know when im like done and need 2 go 2 sleep?

OpenStudy (anonymous):

yes sorry this is so late i need to get this done for my gf

OpenStudy (anonymous):

seriously? she's making u do this??? that's pretty messed up no wait ur pretty messed up that u played the part of the smart guy ur lying to her! :0 uhg! im apalled!

OpenStudy (anonymous):

no i told her i was using help i wouldnt lie to her!!!

OpenStudy (anonymous):

hmmm sure!!!!! thats wat they ALL say even 2 cyber ppl! ugh!!! no its okay i guess just dont lie to her thats mean! okay i'll help up to my limit :)

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

now x=6 do i put that into the equation 4x + 2y = 20?

OpenStudy (anonymous):

bravo! yes sir :)

OpenStudy (anonymous):

y=-2

OpenStudy (anonymous):

no.. y=2

OpenStudy (anonymous):

how 4 times 6 plus 2y equal 20 24 plus 2y equals 20 -24 on both sides then divide by 2

OpenStudy (anonymous):

i'm so sorry i though it was for y my bad :( ur right sorry

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

the next question are you ready?

OpenStudy (anonymous):

sure thang

OpenStudy (anonymous):

\[3x ^{2}y ^{-2}\times2x ^{-3}y ^{2}\div6x ^{2}y ^{2}\times7xy ^{-2}\] they want to simplify it

OpenStudy (anonymous):

this'll take some time i'm solving it ......

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

wait so their both multiplied?

OpenStudy (anonymous):

like individually?

OpenStudy (anonymous):

its hard to do on here 3xsquared times y -squared over 6x squared times y squared times 2x-cubed times ysquared over 7xy -squared

OpenStudy (anonymous):

okay so are 3x^2y^=2 and 2x^x3y^2 multiplied?

OpenStudy (anonymous):

kinda its 3x^2y^-2 over 6x^2y^2 times 2x^-3y^2 over 7xy^-2

OpenStudy (anonymous):

oooooooooooooooooooooooooo okay that makes sense easy!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

after all the cross cancelations u would obtain.... 3x/6x^5 (multiplied by) 2/7 and u when then get 6x/42x^2 simplified would then be x/7x^5

OpenStudy (anonymous):

so x over 7^5

OpenStudy (anonymous):

x over 7x^5

OpenStudy (anonymous):

ok ready for the next one?

OpenStudy (anonymous):

yea can u limit them to a couple now plz?

OpenStudy (anonymous):

2 to 4 ok? i know its going on 1am here

OpenStudy (anonymous):

yeah here it 10:51 pm but i have school tomorow so yaeh its ok. dang ur so determined for ur gurl thats cute! she is very lucky :) keep it going

OpenStudy (anonymous):

hello? u still there?

OpenStudy (anonymous):

c ya im out